OpenStudy (anonymous):

Find the domain and range, and the equation for the axis of symmetry in the following parabola. y=(x-4)²+6

OpenStudy (anonymous):

since \((x-4)^2\geq 0\) on account of it is a square you know \[(x-4)^2+6\geq 6\]

OpenStudy (anonymous):

tells you the range is \([6\infty)\)

OpenStudy (anonymous):

ok

OpenStudy (experimentx):